# and quantum information science

### v2

@@ 8.2 Holographic path-integral optimization
Figure 14: Holographic proposal for the gravity dual of the path-integral optimization, adapted from [421].
@@
With this motivation, references [421, 422] proposed a holographic dual of the path-integral optimization procedure. Their construction was closely related to the framework of Anti-de Sitter/Boundary Conformal Field Theory (AdS/BCFT) [387] in which one considers only a portion of the holographic geometry bounded by an end-of-the-world surface.
@@
Now, the main observation of [421, 422] was that a maximization of the semiclassical HH wave function with respect to the choice of $\phi(\omega)$ is equivalent to imposing the Neumann boundary conditions (in analogy with AdS/BCFT) on $Q$
@@
This agreement was conjectured (interpreted) as the correspondence between the maximization of HH wave functions in AdS and the path integral optimization in holographic CFTs. It is also interesting to comment that parameter $\mu$ in the Liouville action effectively quantifies the amount of optimization, ranging from the unoptimized original metric for $\mu=0$ and the optimal for $\mu=1$. This way, tension $T$ can be seen as its holographic counterpart as it parametrizes a family of surfaces $Q$ starting from the boundary surface $\Sigma$ for $T=-(d-1)$ up to surface $\tau=0$ for $T=0$ (see [421] for detailed discussion).
@@
The second key insight from this construction was a UV completion of the Liouville action, which could be interpreted as specifiying the leading $1/\epsilon$ contribution to path-integral complexity (see the discussion in [420]). Specifically, [421, 422] argued that the gravity action in region $M$ (see Fig. 14), supplemented with appropriate boundary and Hayward corner terms, provides a finite-cutoff completion of the Liouville action. As a consistency check, taking the "boundary" limit of the gravity action reproduced both the Liouville action (146), and its higher-dimensional generalization [39].
@@
A closely related setup was later explored in [423] as an approach to tensor networks in holographic CFTs. The authors demonstrated that the CFT on the slices $Q$ above could be explicitly constructed using $T\bar{T}$ deformation with a coupling that depends on Euclidean time, [^41] This deformation later became the starting point of the “Cauchy-slice holography” [424]. and could be described within finite cut-off holography [425]. Parallel works [41, 43] also argued for $T\bar{T}$-like interpretation of path integral complexity in holographic CFTs. The key distinction between [421, 422] and [423] was the way in which the partition functions of holographic CFTs deformed by $T\bar{T}$ were computed. Namely, in the $T\bar{T}$ context, the on-shell action was evaluated in the bulk region from the interior up to $Q$; whereas in the holographic path-integral optimization context, the action was evaluated in the region $M$ attached to the boundary. Nevertheless, as functionals of the induced metric on $Q$ (i.e. the Weyl factor $\phi(\omega)$), gravitational actions in both computations are the same; therefore, their maximization yields the same metrics $\phi(\omega)$. In other words, their maximization selects a particular slice $z=f(\tau)$ of asymptotically, locally $\mathrm{AdS}$ spacetimes that has constant extrinsic curvature $K$, the so-called constant-mean-curvature (CMC) slices.
@@
Interestingly, slices of AdS spacetime with constant extrinsic curvature also appear naturally in the context of bulk and boundary symplectic forms [388]. Reference [48] proposed the same kind of slices to match the volume of the Einstein-Rosen Bridge in three-dimensional black hole geometries with a boundary notion of Nielsen complexity.
@@ 9 Quantum complexity in quantum information theory and many-body physics
Quantum information (QI) theory and complexity inform each other in many ways. We focus on three. First, QI theory shows that complexity governs a resource consumed in operational tasks (Sec. 9.1). Second, complexity distinguishes topological phases (Sec. 9.2). Third, a QI-theoretic quantity—entanglement entropy—bounds complexity under certain conditions (Sec. 9.3). [^42] We already discussed another connection between complexity and entanglement in the context of Nielsen’s complexity in section 5.5. There, we saw that very special definitions of complexity are completely fixed by the entanglement and Rényi entropies. Finally, the probability distribution over bitstrings $|x\rangle,x\in\{0,1\}^{n}$ induced by measuring a state also serves as a tool for bounding complexity (Sec. 9.4). Other QI-complexity intersections that we do not discuss extensively below include the bounding of complexity in terms of the quantum Wasserstein distance [426], the relationship between complexity and the expressivity of quantum circuits [427], and applications of complexity to quantum machine learning [428].
@@ 9.1 Operationalism and resource theory
The authors of [55] suggested that complexity controls a resource useful in operational tasks. Let $\rho$ denote an arbitrary quantum state of a system $S$ and let $\mathcal{C}_{0}(\rho)$ denote the exact circuit complexity of $\rho$: the least number of basic operations required to prepare $\rho$. Let $\mathcal{C}_{\rm max}$ denote the greatest possible complexity of any state on $S$. For an $N$-qubit system, $\mathcal{C}_{\rm max}\sim 2^{N}\,$. Define the uncomplexity of $\rho$ as the gap $\mathcal{C}_{\rm max}-\mathcal{C}_{0}(\rho)$. Uncomplexity serves as a resource in quantum computation. For example, consider running a quantum computation on an $n$-qubit register. The natural starting point is the zero-complexity state $|0\rangle^{\otimes n}$. It serves analogously to clean scrap paper on which we write scratchwork when performing classical computations.
@@
Brown and Susskind conjectured that one can define a resource theory for quantum uncomplexity [55]. A resource theory models scenarios in which agents are restricted to perform only certain operations. For example, in the resource theory of entanglement, agents can perform only local operations and classical communications (LOCC). The allowed operations are called free, and are modeled as costing nothing. Anything not free is a resource; it may be consumed to enable an operational task. In the entanglement theory, for instance, entangled states are resources. Agents may consume a maximally entangled pair of qubits to teleport quantum information, using LOCC.
@@
A resource theory of quantum uncomplexity was defined in [429]. For simplicity, suppose that an agent in the resource theory has $n$ qubits (the setting can be generalized). The agent can attempt to perform any two-qubit gate $U$ (the setting can be generalized to $k$-qubit gates, for any $k\leq n$, and to gates in a particular set). However, noise corrupts every gate implementation. Hence the gate $\tilde{U}$ that is realized is chosen randomly from an $\epsilon$-ball (a small open set) about $U$. $\tilde{U}$ is called a fuzzy gate. The compositions of all fuzzy gates are the fuzzy operations. Every agent will have a tolerance for how fuzzy, or unknown, the state can become. This tolerance limits how many fuzzy gates the agent is willing to perform—and so curtails the complexities of the operations that will be performed on the system.
@@
Resource theories can be used to formulate operational tasks and to quantify their optimal efficiencies. An example is uncomplexity extraction [429]: let $\delta\geq 0$ denote an error tolerance; and $\rho$, an arbitrary $n$-qubit state. One extracts uncomplexity from $\rho$ by distilling a state $\delta$-close to $|0\rangle^{\otimes k}$ in trace distance. How large a $k$ is achievable? The answer depends on the complexity entropy.
[…]
@@
[…]
Entropies are the workhorses of QI theory. That an entropy can quantify complexity, however, was far from obvious, before [429]. For example, consider an $n$-qubit system prepared in a pure tensor-product state $\rho(0)=\lvert\psi(0)\rangle\!\langle\psi(0)\rvert$. Under a nonintegrable Hamiltonian, the state evolves to $\rho(t)=\lvert\psi(t)\rangle\!\langle\psi(t)\rvert$ in a time $t$. The state’s complexity grows, yet the von Neumann entropy $S_{\rm vN}\bm{(}\rho(t)\bm{)}=-{\rm Tr}\bm{(}\rho(t)\log_{2}\rho(t)\bm{)}$ remains zero. Every subsystem’s reduced state has an $S_{\rm vN}$ that likely grows, but such entropies saturate long before $\rho(t)$’s complexity does. Hence commonly used entropies do not track complexity. Avoiding this pitfall, the complexity entropy (153) quantifies how random $\rho(t)$ appears to a computationally restricted observer. The observer-centric approach was pioneered in [431], through a metric called the strong complexity. Other precursors to the complexity entropy include [432, 433], motivated by pseudorandomness and cryptography. Reference [430] introduces variations on $H_{\rm c}^{r,\eta}(\rho)$. These variations quantify the optimal efficiencies of more information-processing and thermodynamic tasks, including erasure and decoupling.
These results offer hope that complexity entropies might quantify the operational efficiencies of holographic tasks. For example, consider an agent Alice falling into a black hole. A firewall is expected to consume her if the black hole state is dual to a CFT state whose complexity is not increasing [55]. [^43] A subtlety is that, when a black hole’s complexity stops increasing, classical general relativity is expected to stop describing the black hole accurately. Tossing a qubit into the black hole doubles the CFT state’s uncomplexity, enabling the complexity to return to increasing—and saving Alice. This story might be generalized beyond qubits with help from $H_{\rm c}^{r,\eta}$.
@@ 9.2 Circuit complexity and topological order
Until the early 80s it was thought that all phases of matter fit into Landau’s framework of symmetry breaking [434, 435]. Roughly, Landau’s theory asserts that a phase transition occurs when a symmetry is broken, like the spontaneous change from a continuous symmetry in a chaotic/liquid phase to a discrete symmetry in a solid phase. Here, every phase is described by a pair of symmetry groups $G_{H}$ and $G_{\psi}$, where $G_{H}$ are all symmetry operations of the Hamiltonian and $G_{\psi}\subseteq G_{H}$ is the symmetry group of the ground state $\psi$. A phase transition occurs if the symmetry group $G_{\psi}$ changes under an order parameter (such as the strength of a magnetic field).
Starting with a plethora of inequivalent chiral spin states all satisfying the same symmetries, it was discovered that quantum Hamiltonians exhibit an even richer theory of phase transitions. For an extensive review about topological phases, see Wen [436]. Since then an entire “zoo of topological phases” has been discovered [437]. Topological phases refer to gapped ground states that show vast differences in their entanglement structure but are otherwise seemingly featureless (and therefore indistinguishable by symmetry groups).
@@
How does this relate to quantum circuit complexity? It was observed in Ref. [438] that the above definition of topological order of a ground state $|\psi\rangle$ in terms of the absence of adiabatic paths is equivalent to a lower bound that grows in $n$ on the circuit depth for all circuits preparing $|\psi\rangle$ from a product state.
It turns out that that the existence of an adiabatic path is equivalent to the existence of a local unitary transformation [438]:
@@
[…]

### v3

@@ 8.2 Holographic path-integral optimization
Figure 14: Holographic proposal for the gravity dual of the path-integral optimization, adapted from Boruch:2020wax.
@@
With this motivation, references Boruch:2020wax, Boruch:2021hqs proposed a holographic dual of the path-integral optimization procedure. Their construction was closely related to the framework of Anti-de Sitter/Boundary Conformal Field Theory (AdS/BCFT) Takayanagi:2011zk in which one considers only a portion of the holographic geometry bounded by an end-of-the-world surface.
@@
Now, the main observation of Boruch:2020wax, Boruch:2021hqs was that a maximization of the semiclassical HH wave function with respect to the choice of $\phi(\omega)$ is equivalent to imposing the Neumann boundary conditions (in analogy with AdS/BCFT) on $Q$
@@
This agreement was conjectured (interpreted) as the correspondence between the maximization of HH wave functions in AdS and the path integral optimization in holographic CFTs. It is also interesting to comment that parameter $\mu$ in the Liouville action effectively quantifies the amount of optimization, ranging from the unoptimized original metric for $\mu=0$ and the optimal for $\mu=1$. This way, tension $T$ can be seen as its holographic counterpart as it parametrizes a family of surfaces $Q$ starting from the boundary surface $\Sigma$ for T=(d1)T=-(d-1) up to surface $\tau=0$ for $T=0$ (see Boruch:2020wax for detailed discussion).
@@
The second key insight from this construction was a UV completion of the Liouville action, which could be interpreted as specifiying the leading $1/\epsilon$ contribution to path-integral complexity (see the discussion in Camargo:2019isp). Specifically, Boruch:2020wax, Boruch:2021hqs argued that the gravity action in region $M$ (see Fig. 14), supplemented with appropriate boundary and Hayward corner terms, provides a finite-cutoff completion of the Liouville action. As a consistency check, taking the "boundary" limit of the gravity action reproduced both the Liouville action (146), and its higher-dimensional generalization Caputa:2017yrh.
@@
A closely related setup was later explored in Caputa:2020fbc as an approach to tensor networks in holographic CFTs. The authors demonstrated that the CFT on the slices $Q$ above could be explicitly constructed using $T\bar{T}$ deformation with a coupling that depends on Euclidean time, [^44] This deformation later became the starting point of the “Cauchy-slice holography” Araujo-Regado:2022gvw. and could be described within finite cut-off holography McGough:2016lol. Parallel works Chandra:2021kdv, Chandra:2022pgl also argued for $T\bar{T}$-like interpretation of path integral complexity in holographic CFTs. The key distinction between Boruch:2020wax, Boruch:2021hqs and Caputa:2020fbc was the way in which the partition functions of holographic CFTs deformed by $T\bar{T}$ were computed. Namely, in the $T\bar{T}$ context, the on-shell action was evaluated in the bulk region from the interior up to $Q$; whereas in the holographic path-integral optimization context, the action was evaluated in the region $M$ attached to the boundary. Nevertheless, as functionals of the induced metric on $Q$ (i.e. the Weyl factor $\phi(\omega)$), gravitational actions in both computations are the same; therefore, their maximization yields the same metrics $\phi(\omega)$. In other words, their maximization selects a particular slice $z=f(\tau)$ of asymptotically, locally $\mathrm{AdS}$ spacetimes that has constant extrinsic curvature $K$, the so-called constant-mean-curvature (CMC) slices.
@@
Interestingly, slices of AdS spacetime with constant extrinsic curvature also appear naturally in the context of bulk and boundary symplectic forms Belin:2018fxe. Reference Belin:2018bpg proposed the same kind of slices to match the volume of the Einstein-Rosen Bridge in three-dimensional black hole geometries with a boundary notion of Nielsen complexity.
@@ 9 Quantum complexity in quantum information theory and many-body physics
Quantum information (QI) theory and complexity inform each other in many ways. We focus on three. First, QI theory shows that complexity governs a resource consumed in operational tasks (Sec. 9.1). Second, complexity distinguishes topological phases (Sec. 9.2). Third, a QI-theoretic quantity—entanglement entropy—bounds complexity under certain conditions (Sec. 9.3). [^45] We already discussed another connection between complexity and entanglement in the context of Nielsen’s complexity in section 5.5. There, we saw that very special definitions of complexity are completely fixed by the entanglement and Rényi entropies. Finally, the probability distribution over bitstrings $|x\rangle,x\in\{0,1\}^{n}$ induced by measuring a state also serves as a tool for bounding complexity (Sec. 9.4). Other QI-complexity intersections that we do not discuss extensively below include the bounding of complexity in terms of the quantum Wasserstein distance Li_22_Wasserstein, the relationship between complexity and the expressivity of quantum circuits Haug_21_Capacity, and applications of complexity to quantum machine learning Haug_24_Generalization.
@@ 9.1 Operationalism and resource theory
The authors of Brown:2017jil suggested that complexity controls a resource useful in operational tasks. Let $\rho$ denote an arbitrary quantum state of a system $S$ and let $\mathcal{C}_{0}(\rho)$ denote the exact circuit complexity of $\rho$: the least number of basic operations required to prepare $\rho$. Let $\mathcal{C}_{\rm max}$ denote the greatest possible complexity of any state on $S$. For an $N$-qubit system, $\mathcal{C}_{\rm max}\sim 2^{N}\,$. Define the uncomplexity of $\rho$ as the gap $\mathcal{C}_{\rm max}-\mathcal{C}_{0}(\rho)$. Uncomplexity serves as a resource in quantum computation. For example, consider running a quantum computation on an $n$-qubit register. The natural starting point is the zero-complexity state $|0\rangle^{\otimes n}$. It serves analogously to clean scrap paper on which we write scratchwork when performing classical computations.
@@
Brown and Susskind conjectured that one can define a resource theory for quantum uncomplexity Brown:2017jil. A resource theory models scenarios in which agents are restricted to perform only certain operations. For example, in the resource theory of entanglement, agents can perform only local operations and classical communications (LOCC). The allowed operations are called free, and are modeled as costing nothing. Anything not free is a resource; it may be consumed to enable an operational task. In the entanglement theory, for instance, entangled states are resources. Agents may consume a maximally entangled pair of qubits to teleport quantum information, using LOCC.
@@
A resource theory of quantum uncomplexity was defined in YungerHalpern:22Resource. For simplicity, suppose that an agent in the resource theory has $n$ qubits (the setting can be generalized). The agent can attempt to perform any two-qubit gate $U$ (the setting can be generalized to $k$-qubit gates, for any $k\leq n$, and to gates in a particular set). However, noise corrupts every gate implementation. Hence the gate $\tilde{U}$ that is realized is chosen randomly from an $\epsilon$-ball (a small open set) about $U$. $\tilde{U}$ is called a fuzzy gate. The compositions of all fuzzy gates are the fuzzy operations. Every agent will have a tolerance for how fuzzy, or unknown, the state can become. This tolerance limits how many fuzzy gates the agent is willing to perform—and so curtails the complexities of the operations that will be performed on the system.
@@
Resource theories can be used to formulate operational tasks and to quantify their optimal efficiencies. An example is uncomplexity extraction YungerHalpern:22Resource: let $\delta\geq 0$ denote an error tolerance; and $\rho$, an arbitrary $n$-qubit state. One extracts uncomplexity from $\rho$ by distilling a state $\delta$-close to $|0\rangle^{\otimes k}$ in trace distance. How large a $k$ is achievable? The answer depends on the complexity entropy.
The complexity entropy quantifies how random $\rho$ looks to an agent who can perform only computationally restricted measurements YungerHalpern:22Resource, Munson:2024Complexity. In QI theory, a measurement is modeled with a positive-operator-valued measure $\{Q_{\alpha}\}$ nielsen2010quantum. The measurement operators $Q_{\alpha}$ are positive-semidefinite, $Q_{\alpha}\geq 0$, and normalized as $\sum_{\alpha}Q_{\alpha}=\mathbbm{1}$. Define as a zero-complexity measurement operator any $Q_{\alpha}$ that acts on each qubit either (i) as the identity or (ii) by projecting the qubit onto $\lvert 0\rangle\!\langle 0\rvert$: $\bigotimes_{j=1}^{n}(\lvert 0\rangle\!\langle 0\rvert)^{\alpha_{j}}$, if $\alpha_{j}\in\{0,1\}$, and $(\lvert 0\rangle\!\langle 0\rvert)^{0}\coloneqq\mathbbm{1}$. Let $r\in\mathbbm{Z}_{\geq 0}$ denote a fixed integer. Consider acting with $\leq r$ gates and then a zero-complexity measurement operator. This process effects a complexity-$r$ measurement operator, the set of which we denote by $M_{r}$. We apply these definitions as follows. Consider performing a binary hypothesis test between $\rho$ and $\mathbbm{1}/2^{n}$ (the maximally mixed state, and hence the maximally random, state). We receive one of the two states and must guess which we received. We perform a measurement $\{Q,\mathbbm{1}-Q\}$. (In the notation above, $Q=Q_{1}\in M_{r}$, and $\mathbbm{1}-Q=Q_{2}$.) If the $Q$ outcome obtains, we guess that we received $\rho$; and, if the $\mathbbm{1}-Q$ outcome, then $\mathbbm{1}/2^{n}$. A type I error occurs if, upon receiving $\rho$, we wrongfully guess $\mathbbm{1}/2^{n}$. Let $1-\eta$ upper-bound the maximum allowable type I–error probability: ${\rm Tr}(Q\rho)\geq\eta$. A type II error occurs if, upon receiving $\mathbbm{1}/2^{n}$, we wrongfully guess $\rho$. The minimal type II–error probability ${\rm Tr}(Q\mathbbm{1})/2^{n}$, subject to the type I constraint, motivates the complexity entropy:
@@
The complexity entropy controls the number of uncomplex qubits extractable from $\rho$ via fuzzy operations YungerHalpern:22Resource. Let $\epsilon$ quantify the gates’ fuzziness, and let the tolerance $\delta\geq r\epsilon$. For every parameter value $\eta\in[1-(\delta-r\epsilon)^{2},\,1]$, one can extract $k=n-H_{\rm c}^{r,\eta}(\rho)$ qubits $\delta$-close to $|0\rangle^{\otimes n}$, using fuzzy operations. Conversely, one can extract at most $n-H_{\rm c}^{r,1-\delta}$ such qubits. These results follow a pattern common in QI theory: entropies quantify the optimal efficiencies with which operational tasks can be performed.
Entropies are the workhorses of QI theory. That an entropy can quantify complexity, however, was far from obvious, before YungerHalpern:22Resource. For example, consider an $n$-qubit system prepared in a pure tensor-product state $\rho(0)=\lvert\psi(0)\rangle\!\langle\psi(0)\rvert$. Under a nonintegrable Hamiltonian, the state evolves to $\rho(t)=\lvert\psi(t)\rangle\!\langle\psi(t)\rvert$ in a time $t$. The state’s complexity grows, yet the von Neumann entropy $S_{\rm vN}\bm{(}\rho(t)\bm{)}=-{\rm Tr}\bm{(}\rho(t)\log_{2}\rho(t)\bm{)}$ remains zero. Every subsystem’s reduced state has an $S_{\rm vN}$ that likely grows, but such entropies saturate long before $\rho(t)$’s complexity does. Hence commonly used entropies do not track complexity. Avoiding this pitfall, the complexity entropy (153) quantifies how random $\rho(t)$ appears to a computationally restricted observer. The observer-centric approach was pioneered in Brandao:2019sgy, through a metric called the strong complexity. Other precursors to the complexity entropy include Hrastad:1999Pseudorandom, Chen:2017Computational, motivated by pseudorandomness and cryptography. Reference Munson:2024Complexity introduces variations on $H_{\rm c}^{r,\eta}(\rho)$. These variations quantify the optimal efficiencies of more information-processing and thermodynamic tasks, including erasure and decoupling.
These results offer hope that complexity entropies might quantify the operational efficiencies of holographic tasks. For example, consider an agent Alice falling into a black hole. Debate surrounds the possibility that she will encounter a firewall. According to one proposal, a firewall will consume her if the black hole state is dual to a CFT state whose complexity is not increasing Brown:2017jil. [^46] A subtlety is that, when a black hole’s complexity stops increasing, classical general relativity is expected to stop describing the black hole accurately. Tossing a qubit into the black hole doubles the CFT state’s uncomplexity, enabling the complexity to return to increasing—and saving Alice. This story might be generalized beyond qubits with help from $H_{\rm c}^{r,\eta}$.
@@ 9.2 Circuit complexity and topological order
Until the early 80s it was thought that all phases of matter fit into Landau’s framework of symmetry breaking landau1937phys, landau1950symmetry. Roughly, Landau’s theory asserts that a phase transition occurs when a symmetry is broken, like the spontaneous change from a continuous symmetry in a chaotic/liquid phase to a discrete symmetry in a solid phase. Here, every phase is described by a pair of symmetry groups $G_{H}$ and $G_{\psi}$, where $G_{H}$ are all symmetry operations of the Hamiltonian and $G_{\psi}\subseteq G_{H}$ is the symmetry group of the ground state $\psi$. A phase transition occurs if the symmetry group $G_{\psi}$ changes under an order parameter (such as the strength of a magnetic field).
Starting with a plethora of inequivalent chiral spin states all satisfying the same symmetries, it was discovered that quantum Hamiltonians exhibit an even richer theory of phase transitions. For an extensive review about topological phases, see Wen wen2013topological. Since then an entire “zoo of topological phases” has been discovered wen2017colloquium. Topological phases refer to gapped ground states that show vast differences in their entanglement structure but are otherwise seemingly featureless (and therefore indistinguishable by symmetry groups).
@@
How does this relate to quantum circuit complexity? It was observed in Ref. chen2010local that the above definition of topological order of a ground state $|\psi\rangle$ in terms of the absence of adiabatic paths is equivalent to a lower bound that grows in $n$ on the circuit depth for all circuits preparing $|\psi\rangle$ from a product state.
It turns out that that the existence of an adiabatic path is equivalent to the existence of a local unitary transformation chen2010local:
@@
It is easy to see that the Hamiltonians $H(s)$ are all gapped and the unitaries $U(s)$ preserve the spectrum of HH and therefore the spectral gap as well. This is because, if $\lambda$ is an eigenvalue of HH with eigenvector $|\psi\rangle$, then $U(s)|\psi\rangle$ is an eigenvector of $U(s)HU(s)^{\dagger}$ with the same eigenvalue $\lambda$ because $U(s)HU(s)^{\dagger}U(s)|\psi\rangle=U(s)H|\psi\rangle=\lambda U(s)|\psi\rangle$. The relation to circuit complexity now follows from the fact that the operator $U(1)$ can be approximated by a constant depth quantum circuit via Trotterization. Trotterization proceeds by splitting the time evolution (here the integral $\int_{0}^{1}ds\,\tilde{H}(s)$) into small pieces, which are approximately local. The error this approximation introduces can be made arbitrarily small by increasing the circuit depth. In addition to this general result, different notions of quantum complexity presented in this review, including Nielsen and Krylov complexities, have also been explored in examples of specific systems exhibiting topological phase transitions, see, e.g., Liu:2019aji, Xiong:2019xoh, Ali:2018aon, Caputa:2022eye.